OpenStudy (anonymous):

please tell me how to use symbol code, in laTex .......i tried but i cann't do

OpenStudy (anonymous):

@somy @mathstudent55 plz help me

OpenStudy (abhisar):

What symbol codes do you want to use ?

OpenStudy (abhisar):

\(\alpha\\ \beta\\ \gamma\\ \delta\\ \epsilon\\ \zeta\\ \eta\\ \theta\\ \iota\\ \kappa\\ \lambda\\ \mu\\ \nu\\ \xi\\ \pi\\ \rho\\ \sigma\\ \tau\\ \upsilon\\ \phi\\ \chi\\ \psi\\ \omega\\ \Gamma\\ \Delta\\ \Theta\\ \Lambda\\ \Xi\\ \Pi\\ \Sigma\\ \Gamma\\ \Phi\\ \Psi\\ \Omega\\ \infty\)

OpenStudy (abhisar):

@ziqbal103 let me know if you are looking for more symbols.

OpenStudy (abhisar):

`\(\alpha\\ \beta\\ \gamma\\ \delta\\ \epsilon\\ \zeta\\ \eta\\ \theta\\ \iota\\ \kappa\\ \lambda\\ \mu\\ \nu\\ \xi\\ \pi\\ \rho\\ \sigma\\ \tau\\ \upsilon\\ \phi\\ \chi\\ \psi\\ \omega\\ \Gamma\\ \Delta\\ \Theta\\ \Lambda\\ \Xi\\ \Pi\\ \Sigma\\ \Gamma\\ \Phi\\ \Psi\\ \Omega\\ \infty\)` Two codes are seperated by \\

OpenStudy (anonymous):

Syntax for symbols: ∖(∖symbolcode∖) symbol←⇐⟵⟸↔⟷↑≤≪≈±∑♢△LATEXcode∖leftarrow∖Leftarrow∖longleftarrow∖Longleftarrow∖leftrightarrow∖longleftrightarrow∖uparrow∖leq∖ll∖approx∖pm∖sum∖diamondsuit∖triangle∖LaTeXsymbol→⇒⟶⟹⇔⟺↓≥≫≠÷∞♡□★code∖righttarrow∖Righttarrow∖longrighttarrow∖Longrighttarrow∖Leftrightarrow∖Longleftrightarrow∖downarrow∖geq∖gg∖neq∖div∖infty∖heartsuit∖Box∖bigstar Example of combination with size and color:

OpenStudy (anonymous):

@abhisar these heart suit ,stars and otthers that are above

OpenStudy (anonymous):

\(\heartsuit\)

OpenStudy (anonymous):

\(\large\red\heartsuit\)

OpenStudy (anonymous):

\(\red\bigstar\)

OpenStudy (anonymous):

\(red\huge{bigstar}\)

OpenStudy (anonymous):

\(.~~.\) \(\smile\)

OpenStudy (anonymous):

\(\SCR\{red}{i~am~fine}\) Now this italic text is nice, but we can use way cooler fonts! The syntax for this is: ∖(∖font text∖) ∖(∖bf Hello∼there!∖) Will become: Hello there! Fonts you can use: ∖sf→ABCDEFabcdef123456 ∖tt→ABCDEFabcdef123456 ∖it→ABCDEFabcdef123456 ∖bf→ABCDEFabcdef123456 ∖rm→ABCDEFabcdef123456 ∖cal→ABCDEFabcdef123456 ∖scr→ABCDEFabcdef123456 ∖frak→ABCDEFabcdef123456 There are 2 exceptions: You need to add braces to these fonts in order to use them, so ∖(∖mathbb{text}∖) ∖mathbb→ABCDEFabcdef123456 ∖boldsymbol→ABCDEFabcdef123456

OpenStudy (anonymous):

@Abhisar and these as well \(\scr\rightarrow\{ABCD\)

OpenStudy (anonymous):

\(\FRAC\{ABCD\)

OpenStudy (anonymous):

\(\Huge\color{\red}{\rightarrow}\)

OpenStudy (anonymous):

\(\huge\leftarrow\) \(\green\Huge\downarrow\)~\(\color\purple\Huge\{abovearrow}\)

OpenStudy (anonymous):

@abhisar there is mistake above

OpenStudy (anonymous):

\(\mathbb{WELCOME~ON~OPEN~STUDY}\)

OpenStudy (anonymous):

\(\Huge\color{green}{I~am~trying~to~do~good~''LaTeX''~\color {purple}Abhisar}\)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @ziqbal103 \(\huge\leftarrow\) \(\color{green}{\Huge\downarrow}\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

``` \(\huge\leftarrow\) \(\color{green}{\Huge\downarrow}\) ```

OpenStudy (anonymous):

i want to write text in huge purple size in(\{scr}\)@Abhisar as i am trying downn \(\scr{color}{purple}~{WELCOME HERE}\) HERE color purple is not showing @Abhisar

OpenStudy (anonymous):

\(\HUGE\{color}{pink}{uparrow\) what is the mistake here?@Abhisar

OpenStudy (anonymous):

\(\Huge\color{pink}{\uparrow}\)\(\Huge\color{brown}{\downarrow}\)\(\Huge\color{purple}{\rightarrow}\)\(\Huge\color{grey}{\leftarrow}\)

OpenStudy (anonymous):

\(\Huge\color{blue}{Abhisar~in~which~CLASS~u~r?}\)

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

\(\cal\Huge\color{orange}{reply~me~fast}\)

OpenStudy (anonymous):

\(\bf\Huge\color{brown}{i~am~trying~to~learn~latex?}\)

OpenStudy (anonymous):

how to do big\(\huge\heartsuit\)

OpenStudy (anonymous):

@abhisar \(THANKS\)